OpenStudy (anonymous):

WHAT IS THE X INTERCEPT OF 2X^2+3X-20=0

OpenStudy (anonymous):

2x^2 +8x -5x -20 = 0 (x+4)(2x-5) = 0

OpenStudy (anonymous):

so x = -5 or 5/2

OpenStudy (anonymous):

Wat @Kanwar245 just said is how to find the y intercept unless that is wat u mean. But to find the x intersept all u do is sub in x=0 and solve for y

OpenStudy (anonymous):

are u sure @HawkCrimson

OpenStudy (anonymous):

yes and ill draw a graph to prove it

OpenStudy (anonymous):

oh dont worry

OpenStudy (anonymous):

lol! HE WILL GRAPH THAT UP

OpenStudy (anonymous):

wups im wrong lol

OpenStudy (anonymous):

 that there is the two points where the curve cuts the X axis which is wat u found, my bad

OpenStudy (rc94):

Kanwar was right except he meant -4 not -5 haha

OpenStudy (anonymous):

yep